Euler’s constant

γ = lim_{n→∞} (\frac{1}{1} + \frac{1}{2} + ⋯ + \frac{1}{n} − log n)

1962 Knuth  1271  Euler-MacLaurin
1963 Sweeney  3683 20 h exponential integral
1977 Brent  20700 20 h exponential integral
1980 Brent, McMillan 30100  Bessel integral
1997 Papanikolaou  500000 160 h Bessel integral
1997 Papanikolaou 1000000 42 h Bessel integral, bin. splitting